# One-Token Rollout: Guiding Supervised Fine-Tuning of LLMs with Policy Gradient

### v1

@@ 6 Conclusion
In this work, we investigated the generalization weakness of SFT compared to RL, positing that the disparity stems from the fundamental difference between SFT’s static, off-policy data and RL’s dynamic, on-policy data. To bridge this gap from a data-centric perspective, we introduced One-Token Rollout, a novel fine-tuning algorithm. By reframing each token generation as a single-step reinforcement learning trajectory, OTR transforms the static supervised dataset into a dynamic, on-policy learning signal, successfully incorporating the advantage of on-policy data into the SFT framework while maintaining its computational efficiency. Our extensive experiments empirically validate this approach, demonstrating that OTR consistently outperforms SFT on a wide array of in-domain and out-of-domain benchmarks. Ultimately, we present OTR as a powerful and practical alternative for fine-tuning LLMs, providing compelling evidence that simulating on-policy interaction is a key direction for developing more generalizable fine-tuned language models.
Appendix A Additional Experiment
@@
Limitations

### v2

@@ 6 Conclusion
In this work, we investigated the generalization weakness of SFT compared to RL, positing that the disparity stems from the fundamental difference between SFT’s static, off-policy data and RL’s dynamic, on-policy data. To bridge this gap from a data-centric perspective, we introduced One-Token Rollout, a novel fine-tuning algorithm. By reframing each token generation as a single-step reinforcement learning trajectory, OTR transforms the static supervised dataset into a dynamic, on-policy learning signal, successfully incorporating the advantage of on-policy data into the SFT framework while maintaining its computational efficiency. Our extensive experiments empirically validate this approach, demonstrating OTR consistently outperforms SFT on a wide array of in-domain and OOD benchmarks. Ultimately, we present OTR as a powerful and practical alternative for fine-tuning LLMs, providing compelling evidence that simulating on-policy interaction is a key direction for developing more generalizable fine-tuned language models.
Limitations
Appendix A Computational Efficiency Analysis
To assess the computational overhead of our proposed method, we conduct a comprehensive comparison of training efficiency between OTR and the SFT baseline. We report the total training time and peak GPU memory usage (VRAM) across three different model scales: Qwen2.5-7B, Qwen3-8B, and Olmo3-7B.
Table 5 presents the quantitative results. As evidenced by the data, OTR incurs only a slight increase in training time compared to SFT. This is primarily because OTR performs rollout for only a single token step, avoiding the expensive generation costs typically associated with standard RL methods that require full-sentence rollouts. In the case of Qwen3-8B, OTR even recorded a slightly shorter training time; we attribute such minor discrepancies to normal system fluctuations and data loading variances rather than algorithmic differences. In terms of memory consumption, OTR exhibits a marginal increase in peak VRAM usage compared to SFT. A key factor contributing to this memory efficiency is that the sampling process is wrapped in a torch.no_grad() context. Consequently, SFT serves as a fair and strong baseline for evaluating the performance gains of OTR under a strict computational budget.
Appendix B Additional Experiment
@@
Table 5: Comparison of computational efficiency between SFT and OTR.
Model Method Training Time Peak Memory (GB)
Qwen2.5-7B SFT 2h 06m 72.88
OTR 2h 07m 75.89
Qwen3-8B SFT 2h 39m 76.97
OTR 2h 30m 79.60
Olmo3-7B SFT 2h 21m 71.00
OTR 2h 22m 72.92
@@
Appendix C Comparison with Dynamic Fine-Tuning
The OTR framework can be seen as a generalization of DFT. This relationship becomes clear when we consider the special case of our OTR objective where $N_{gt}\neq 0$ and the hyperparameter $\beta=0$. In this scenario, the second term in Equation 9, which penalizes incorrect samples, vanishes. Then the OTR loss can be formulated as:
$\mathcal{L}_{\text{OTR}}^{t}(\theta)|_{\beta=0}=-\frac{N_{gt}}{K}\log\pi_{\theta}(x_{t}|s_{t}).$ (12)
where $N_{gt}/K$ represents the empirical frequency of sampling the ground-truth token during the rollout. This frequency is a direct Monte Carlo approximation of the ground-truth token’s probability, i.e., $\frac{N_{gt}}{K}\approx\pi_{\theta}(x_{t}|s_{t})$. Thus, when $N_{gt}\neq 0$ and $\beta=0$, the OTR objective is functionally equivalent to the DFT objective, as both methods effectively weight the loss of the ground-truth token by its estimated probability.